OpenStudy (anonymous):

Change the equation to slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all your work

jagr2713 (jagr2713):

The equation ?

OpenStudy (anonymous):

Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold.

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@jagr2713

OpenStudy (alexandervonhumboldt2):

to change the equation to slope intercept form (y=mx+b) you first need to subtract 2x from both sides: 2x + 3y = 1,470 2x-2x+3y=1470-2x 3y=1470-2x now divide both sides by 3: 3y/3=(1470-2x)/3 y=-2/3x + 490

OpenStudy (alexandervonhumboldt2):

in the y=mx+b form m is slope and b is y intercept

OpenStudy (alexandervonhumboldt2):

y=mx+b y=-2/3x +490 what do you think the y-inter and slope is in your case?

OpenStudy (alexandervonhumboldt2):

remember: slope in slope intercept form y=mx+b is m slope in standard form ax+by=c is -a/b